you just bought an investment property for $450,000 . The loan is fully amortized over 30 years and payments are made monthly You took a loan for 80% of the purchase price to buy the property. The nominal annual rate of interest is 3.9% for this loan. a. How much will your monthly loan payments be? b. How much total interest will you pay over the entire life of the loan? What is your loan balance after exactly 15 years of payments (180 payments) have been made? c.
Answer #1

loan amount

=80%*450000

=360000

monthly payments=360000/((1-(1+(3.9%/12))^(-30*12))/(3.9%/12))=1698.01 is answer

total interest to pay over the term of loan=1698.01*360-360000=251283.60 is answer

loan balance after 15 years=1698.01*((1-(1+(3.9%/12))^(-15*12))/(3.9%/12))=231120.29 is answer

the above are answers